Operation Summary for April 28, 2010 - April 30, 2010 |
Wednesday April 28 The day shift began with Operations monitoring the Tevatron (TeV) store 7771 (304.0E30), the Recycler stash (51.0E10), and the Antiproton Source stack (16.0E10), with the Linac, Booster, and Main Injector operating normally, with the NuMI, MiniBooNE, and MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations reported a problem with the Neutron Therapy Facility 32˚ bend magnet power supply; a power supply person investigated and fixed the problem. FESS personnel tested the MINOS #3 sump pump. Operators prepared the Pelletron for a one-hour access. During the evening shift, Operations disabled the MTest beam at 6 PM. Operations terminated store 7771 at 9:10 PM. Operators prepared D0 for an access. Operations put the TeV into shot setup at 11:42 PM. Operations reported minor downtime due to Booster RF trips and to Pbar lithium lens power supply trips. Thursday April 29 The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (392.9E10), with Pbar stacking (21.1E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established store 7774 at 1:01 AM with an average initial luminosity (AIL) of 350.4E30. Operations established beam to MTest at 4:30 AM. During the day shift, Operations reported some minor downtime due to trips of Linac, Booster, and Pbar Accumulator RF stations (LRF1, BRF7, and ARF1), due to the trip of a Linac RF Buncher, and due to a Meson Lambertson (MLAM) not ramping. MTA ran a beam test and took 60 pulses of beam between 10:09 AM and 12:35 PM. During the evening shift, Operations terminated store 7774 at 4:47 PM, put the TeV into shot setup, but the setup was delayed considerably by TeV RF problems, most of which were traced to an IP address conflict between the operational system and the development system. Operations disabled the MTest beam at 4:58 PM. Shot setup resumed at 10:04 PM and Operations established store 7778 at 11:12 PM with an AIL of 378.3E30. Friday April 30 The midnight shift began with Operations monitoring store 7778 (277.51E30) and the stash (99.7E10), with Pbar stacking (28.0E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, LRF5 tripped off; it reset. Operations established beam to MTest at 4:12 AM. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. After Operations terminates store 7778 this afternoon, the Run Coordinator will allow the following jobs before the next store goes in: Pbar personnel will search for an LCW leak, CDF personnel will conduct detector work, cryo system personnel will conduct some wet engine maintenance, and FESS will test the newly replaced pond water pump at MI-10. |
|
More Information Comments and Suggestions |
| last modified 04/30/2010 email Fermilab |
| Security, Privacy, Legal | |